OpenStudy (anonymous):

which line is not perpendicular to the line 4x-3y=2

OpenStudy (anonymous):

A lot of liens are not perpendicular to 4x-3y=2.

OpenStudy (anonymous):

3x + 4y = 2 3x − 4y = 2 -3x − 4y = 5 3x + 4y = -8

OpenStudy (anonymous):

the suare boxes are both minus sighns

OpenStudy (anonymous):

square*

OpenStudy (anonymous):

3x + 4y = -8 and -3x − 4y = 5 are both not perpendicular to 4x-3y=2

OpenStudy (midhun.madhu1987):

Perpendicular lines have the product of their slopes equal to -1.